OpenStudy (anonymous):

cos(x-π/6)=-.5 Find the exact solutions between 0 ≤ x ≤2π

OpenStudy (anonymous):

\[ \cos( x -\frac \pi 6)= \cos(2 \frac \pi 3) \]

OpenStudy (anonymous):

\[ x -\frac \pi 6= \pm \frac {2 \pi} 3 + 2 k \pi \]

OpenStudy (anonymous):

\[ x = \frac \pi 6\pm \frac {2 \pi} 3 + 2 k \pi \] Choose the ones between 0 and 2 pi and you are done

OpenStudy (anonymous):

i don't understand what you just did there

OpenStudy (anonymous):

the last part i mean

OpenStudy (anonymous):

i mean how do i pick one?

OpenStudy (anonymous):

\[ \cos(u) = \cos(w) \implies x =\pm w + 2 k \pi \]

OpenStudy (anonymous):

You must get two solutions

OpenStudy (anonymous):

\[ x = \frac \pi 6+ \frac {2 \pi} 3 + 2 0 \pi= \frac{ 5\pi} 6 \] That is one

OpenStudy (anonymous):

oh how did you know it was plus or minus 2π/3

OpenStudy (anonymous):

It is not 20 pi it is 2 (0) pi=0 im the above post

OpenStudy (anonymous):

\[ \cos \left(\pm\frac{2 \pi }{3}\right)=-\frac{1}{2} \]

OpenStudy (anonymous):

is it easier to graph it?

OpenStudy (anonymous):

\[\cos \left( x-\frac{ \pi }{6 } \right)=-\frac{ 1 }{ 2 }=-\cos \frac{ \pi }{ 3 }=\cos \left( \pi-\frac{ \pi }{ 3 } \right) ,\cos \left( \pi+\frac{ \pi }{3 } \right)\] \[\cos (x-\frac{ \pi }{6})=\cos \frac{ 2\pi }{3 },\cos \frac{ 4\pi }{3 }\] \[x-\frac{ \pi }{ 6 }=\frac{ 4\pi }{6 }+2k \pi,\frac{ 8\pi }{6 }+2k \pi ,where~ k ~is~ an ~integer.\] \[x=\frac{ 5\pi }{6 }+2k \pi,\frac{ \ 3\pi }{2 }+2k \pi ,where~ k~ is~ an~ integer\]~

OpenStudy (anonymous):

could you draw out what you did there?

OpenStudy (anonymous):

Here is the graph

OpenStudy (anonymous):

how do i turn the decimals so it looks like pi/3 or something? i don't quite understand how to solve for x. am i suppose to add the .5 to the other side of the equation?

OpenStudy (anonymous):

Please Read my older posts. I gave you some details

OpenStudy (anonymous):

oh yeah i saw that, i didn't get how you got 2pi/3

OpenStudy (anonymous):

I told you why \[ \cos( 2 \pi/3)=-\frac 1 2=-.5 \]

OpenStudy (anonymous):

and \[ \cos( -2 \pi/3)=-\frac 1 2=-.5 \]

OpenStudy (anonymous):



OpenStudy (anonymous):